# Faster Motion Planning via Restarts

### v1

@@ 4.1 Counter search strategy
Here we describe a strategy due to Luby et al. [LSZ93] – our exposition is different and is included here for the sake completeness.
@@ 4.1.1 $k$-front
Let alg be a Las-Vegas algorithm. The work quantity associated with alg can be visualized on a 2D system of axes as the area of an axis aligned rectangle with corners at the origin and $\mathrm{profile}\left(\texttt{alg}\right)$ (viewed as a 2D point). Thus, all the algorithms with the same amount of work $\alpha$, have their profile points on the hyperbola $\mathcalb{h}(\alpha)\equiv(y=\alpha/x)$. See Figure 4.1 for an illustration.
Figure 4.1: An illustration of the work associated with four different Las-Vegas algorithms with different profiles $(1/\mathcalb{p},\mathcalb{t})$. The upper right corner of every rectangle is its profile. $\mathcalb{p}$ and $\mathcalb{t}$ are explicitly given in the legend.
A natural way to figure out all the profiles of algorithms that the prefix $t_{1},\ldots,t_{k}$ (of a sequence $\mathcal{T}$), suffices for in expectation ,i.e., a strategy for alg should have a successful run using only these TTLs), is to sort them in decreasing order, and interpret the resulting numbers $t_{1}^{\prime}\geq t^{\prime}_{2},\ldots,t_{k}^{\prime}$ as a bar graph (each bar having width $1$). All the profiles under this graph, are “satisfied” in expectation using this prefix. We refer to the function formed by the top of this histogram, as the $k$-front of $\mathcal{T}$. See Figure 4.2 for an example.
Figure 4.2: An example of an $8$-front of a sequence of TTLs $\mathcal{T}=[7,7,3,2,2,1,1,1...]$. Three parabolas for algorithms with different associated work quantities are plotted alongside the $8$-front of $\mathcal{T}$. Two profiles of algorithms that, in expectation, may not have a successful run are labeled. Note that since we assume that the minimum runtime of an algorithm is one second, the part of the plot where $x<1$ is inconsequential, and the intersection of the parabola with the vertical line $x=1$ is relevant only if the algorithm is deterministic (meaning $\mathcalb{p}=1$).
A key insight is that even if we knew that the profile of alg lies on the hyperbola $\mathcalb{h}(\alpha)\equiv(y=\alpha/x)$, we do not know where it lies on this hyperbola. Thus, an optimal sequence front should sweep over the $\alpha$-work hyperbola
@@
more or less in the same time.
@@ 4.1.2 The strategy
In the following, let $\{x\}^{y}$ be the sequence of length $y$ made out of $x$s. Fix a value $2^{i}$. We are trying to “approximate” the hyperbola $\mathcalb{h}(2^{i})$ with a sequence that its front roughly coincides with this hyperbola. One way to do so, is to consider a sequence $S_{1}$ containing the numbers $\{2^{i}\}^{1},\{2^{i-1}\}^{2},\ldots\{2^{0}\}^{i}$ (not necessarily in this order!).
The intuition for this choice is the following – we might as well round the optimal threshold of the algorithm to the closest (bigger) power of $2$, and let $\nabla$ denote this number. Unfortunately, even if we knew that the work of the optimal simulation is $\alpha=2^{i}$, we do not know which value of $\nabla$ to use, so we try all possible choices for this value. It is not hard to verify that $S_{i}$ dominates the work hyperbola $\mathcalb{h}{2^{i}}$ (for all integral points).
Figure 4.3: Plots of three $k$-fronts of TTL sequences generated by the counter strategy. Note that the bars “hide” each other, e.g. the bars of $S_{3}$ (light purple) have $y$-values (8, 4, 4, 2,…). In particular, the region covered by $S_{1}=2,1,1$ is bounded by the blue staircase. Similarly, $S_{2}=4,2,2,1,1,1,1$ is bounded by the red staircase, and $S_{3}=8,4,4,2,2,2,2,1,\ldots,1$ is bounded the purple staircase.
What is not clear is how to generate an infinite sequence that its prefix is $S_{i}$ if one reads the first $|S_{i}|$ elements of this sequence, for all $i$. As Luby et al. [LSZ93] suggest, ignoring the internal ordering of $S_{i}$, we have that $S_{i+1}=S_{i},S_{i},2^{i+1}$, where $S_{0}=1$. Indeed, every number in $S_{i}$ appears twice in $S_{i+1}$ except for the largest value, which is unique. We have $S_{1}=1,1,2$, $S_{2}=1,1,2,1,1,2,4$, and so on. See Figure 4.3 for an illustration of $k$-fronts generated by the strategy.
Algorithmically, there is a neat way to generate this counter strategy TTL sequence. Let $c$ be an integer counter initialized to 00. To generate the next elements in the sequence, at any point in time, we increase the value of $c$ by one, and add all the numbers $2^{i}$, for $i=0,1,2,3,\ldots$, that divides $c$ into the sequence. Thus, the sequence is
@@
Let $S(i,k)$ be the sum of all the appearances of $2^{i}$ in first $k$ elements of $T_{k}$. For $2^{i}>2^{j}$ (both appearing in the first $k$ elements), we have $S(i,k)\leq S(j,k)\leq 2S(i,k)$, as one can easily verify.
Let $(1/\beta,\Delta)$ be the profile of alg, see Definition 3.5. The optimal simulation takes in expectation $\mathcal{O}=\Theta(\Delta/\beta)$ time. Let $k$ be the number of elements of $T$ that were used by the simulation before it stopped. Let $B$ be the number of elements in $t_{1},\ldots,t_{k}$, at least as large as $\Delta$. Clearly, $B\sim\mathrm{Geom}(\beta)$ (to be more precise it is stochastically dominated by it). Thus, we have ${\mathbb{E}}\left[B\right]\leq 1/\beta$. Furthermore, the quantity $\ell=1+\left\lceil{\log_{2}(B\Delta)}\right\rceil$ is an upper bound on the largest index $\ell$ such that $2^{\ell}$ appears in $t_{1},\ldots,t_{k}$. An easy calculation shows that ${\mathbb{E}}\left[\ell\right]=O(1+\log(\Delta/\beta))=O(\log\mathcalb{w}_{\texttt{alg}})$. The total work of the simulation thus is $\sum_{i=0}^{\ell}S(i,k)=O(\ell B\Delta)$, which in expectation is $O(\mathcalb{w}_{\texttt{alg}}\log\mathcalb{w}_{\texttt{alg}})=O(\mathcal{O}\log\mathcal{O})$.
4.2 New strategies
In the following we present and discuss several restarting strategies. We start by introducing randomized TTL strategies, namely random $\zeta$ search and random counter search, in Section 4.3. These strategies randomly generate TTLs whenever a new run starts instead of using a predetermined sequence of TTLs. We prove that both strategies are optimal.
We then introduce two more strategies for which we abandon the “stop/start only” constraint, and allow the use of pausing. The wide search strategy, discussed in Section 4.4, simulates different copies of alg running at different speeds by pausing some runs while allowing others to progress. The cache + X strategy, which is, in fact, the combination of a caching mechanism and a restarting strategy, maintains a set of paused runs of alg that are “extended” to longer runs instead of simulating alg from scratch at every step. We provide a proof of optimality only for wide search, as the caching mechanism clearly maintains the optimality of the X component of cache + X.
4.3 The random search
@@
Intuitively, our goal is to generate a distribution $\mathcal{D}$ such that the $k$-front created by sampling $\mathcal{D}$ $k$ times will approximate the space under the hyperbola $\mathcalb{h}(2^{\log k-1})$. These distributions are asymptotically optimal, and generate thresholds other than $2^{i}$.
4.3.1 The Zeta $2$ distribution
@@
Let $(1/\beta,\Delta)$ be the profile of alg, and let $L=c\cdot\mathcal{O}^{c}$, where $\mathcal{O}=\Theta(\Delta/\beta)$ and $c$ is some sufficiently large constant, let $\mathcal{T}_{L}=t_{1},t_{2},\ldots,t_{L}\sim\zeta(2)$ be the random sequence of TTLs used by the simulation, and let $X\sim\zeta(2)$. We say that a value $t_{i}\in\mathcal{T}_{L}$ is good if $t_{i}\leq L^{2}$. Clearly, the probability that any of the values in $\mathcal{T}_{L}$ is not good is at most $L{\mathbb{P}}\left[X\geq L^{2}\right]=O(1/L)$, since for $i>1$, we have
${\mathbb{P}}\left[X\geq i\right]=\sum_{\ell=i}^{\infty}\frac{6/\pi^{2}}{\ell^{2}}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
$\leq\frac{6}{\pi^{2}}\sum_{\ell=i}^{\infty}\frac{1}{\ell(\ell-1)}=\frac{6}{\pi^{2}}\sum_{\ell=i}^{\infty}\left(\frac{1}{\ell-1}-\frac{1}{\ell}\right)=\frac{6}{\pi^{2}(i-1)}.$
@@
$\tau={\mathbb{E}}\!\left[X\;\middle|\;X\leq L^{2}\right]\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
$=\frac{1}{{\mathbb{P}}\left[X\leq L^{2}\right]}\sum_{i=1}^{L^{2}}\frac{6i}{\pi^{2}i^{2}}=O(\log L)=O(\log\mathcal{O}).$
@@ 4.3.2 Random counter search
Consider a generated threshold $t$ as a binary string $b_{1}b_{2}\ldots,b_{k}$, where the numerical value of this string, is the value it encode in base two: $t=\sum_{i=1}^{k}b_{i}2^{k-i}$.
@@
We generate a random string as follows. Let $s_{1}=1$. Next, in each step, the string is finalized with probability $1/2$. Otherwise, a random bit is appended to the binary string, where the bit is randomly chosen between 00 or $1$ with equal probability. We denote the resulting distribution on the natural numbers BIN. Observe that for $R\sim\textbf{BIN}$, the probability that $R$ length is $k$ (i.e., $\mathcalb{b}\left(R\right)=k$) is $1/2^{k}$, for $k\geq 1$. Also, observe that $R$ has uniform distribution over all the numbers of the same length.
Note that if we were to “round” every value drawn from the BIN down to the nearest value of the form $2^{i}$, we would get the geometric distribution perfectly mimicking the counter search strategy. We refer to the random strategy using BIN to sampling the TTLs as the Random Counter simulation.
@@
Using the notation of Lemma 4.2 we let $(1/\beta,\Delta)$ be the profile of alg, $L=c\cdot\mathcal{O}^{c}$, where $\mathcal{O}=\Delta/\beta$ and $c$ is some sufficiently large constant, and $T_{L}=t_{1},t_{2},\ldots,t_{L}$ be the random sequence used by the simulation. We also denote values $t_{i}\in\mathcal{T}_{L}$ as good if $t_{i}\leq L^{2}$. Again, we get that the probability that $\mathcal{T}_{L}$ contains a value which is not good is bounded by $1/L$. From this point on, we assume all the values of $T_{L}$ are good, and denote this event by $\mathcal{G}$.
@@
$p={\mathbb{P}}\left[X_{i}=1\right]\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad$
$=\mathop{{\mathbb{P}}}\!\left[\texttt{alg}\text{ successfully runs with {TTL}~{}}t_{i}\;\middle|\;t_{i}\geq\Delta\right]{\mathbb{P}}\left[t_{i}\geq\Delta\right]$
@@
The number of values of $T_{L}$ the simulation needs to use until it succeeds is a geometric variable with distribution $\mathrm{Geom}(1/\mathcal{O})$. (Clearly, the probability the algorithm reads all the values of $T_{L}$ diminishes exponentially, and can be ignored.) Thus, the algorithm in expectation reads $1/p=O(\mathcal{O})$ values of $T_{L}$, and each such value has expectation $O(\log\mathcal{O})$. Thus, conditioned on $\mathcal{G}$, the expected running of the simulation is $O(\mathcal{O}\log\mathcal{O})$.
4.4 The wide search
We now remove the constraint allowing us to only start and terminate runs of alg, and introduce the pause/resume operation, allowing us to freeze a run mid execution, and later resume it with no effect on the performed computations. Note that we still refer to all non fixed-TTL strategies as “restarting strategies” for the sake of brevity. The wide search strategy uses pause/resume operations to simulate running many copies of the algorithm in parallel, but in different speeds. Specifically, the $i$th copy of alg, denoted by $A_{i}$, is ran at speed $\alpha_{i}=1/i$, for $i=1,\ldots$. So, consider the simulation immediately after it ran $A_{1}$ for $t$ seconds — here $t$ is the current time of the simulation. During that time the $i$th algorithm $A_{i}$ ran $n_{i}(t)=\left\lfloor{t/i}\right\rfloor$ seconds. Naturally, the simulation instantiates $A_{i}$ at time $i$, since $n_{i}(i)=1$, and then runs it for one second. More generally, the simulation allocates a running time of one second to a copy of the algorithm $A_{j}$ at the time $t$, when $n_{j}(t)$ had increased to a new integral value. Then, $A_{j}$ is resumed for one second. This simulation scheme can easily be implemented using a heap, and we skip the low-level details.

### v2

@@ 3.1 Counter search strategy
Here we describe a strategy due to Luby et al. [LSZ93] – our exposition is different and is included here for the sake of completeness.
@@ 3.1.1 $k$-front
Let alg be a Las Vegas algorithm. The work quantity associated with alg can be visualized on a 2D system of axes as the area of an axis-aligned rectangle with corners at the origin and $\mathrm{profile}\left(\texttt{alg}\right)=(x,y)$ (viewed as a 2D point) – see Eq. (2.1). Thus, all the algorithms with the same amount of work $\alpha$ have their profile points on the hyperbola $\mathcalb{h}\left(\alpha\right)\equiv(y=\alpha/x)$. See Figure 3.1 for an illustration.
Figure 3.1: An illustration of the work associated with four different Las Vegas algorithms with different profiles $(1/\mathcalb{p},\mathcalb{t})$. The upper right corner of every rectangle is its profile. $\mathcalb{p}$ and $\mathcalb{t}$ are explicitly given in the legend.
A natural way to figure out all the profiles of algorithms that the prefix $t_{1},\ldots,t_{k}$ (of a sequence $\mathcal{T}$), suffices for in expectation, i.e., a strategy for alg should have a successful run using only these TTLs), is to sort them in decreasing order, and interpret the resulting numbers $t_{1}^{\prime}\geq t^{\prime}_{2},\ldots,t_{k}^{\prime}$ as a bar graph (each bar having width 111). All the profiles under this graph are “satisfied” in expectation using this prefix. We refer to the function formed by the top of this histogram as the $k$-front of $\mathcal{T}$. See Figure 3.2 for an example.
Figure 3.2: An example of an 888-front of a sequence of TTLs $\mathcal{T}=[7,7,3,2,2,1,1,1...]$. Three parabolas for algorithms with different associated work quantities are plotted alongside the 888-front of $\mathcal{T}$. Two profiles of algorithms that, in expectation, may not have a successful run are labeled. Note that since we assume that the minimum runtime of an algorithm is one second, the part of the plot where $x<1$ is inconsequential, and the intersection of the parabola with the vertical line $x=1$ is relevant only if the algorithm is deterministic (meaning $\mathcalb{p}=1$).
A key insight is that even if we knew that the profile of alg lies on the hyperbola $\mathcalb{h}\left(\alpha\right)\equiv(y=\alpha/x)$, we do not know where it lies on this hyperbola. Thus, an optimal sequence front should sweep over the $\alpha$-work hyperbola
@@
more or less at the same time.
@@ 3.1.2 The strategy
In the following, let $\{x\}^{y}$ be the sequence of length $y$ made out of $x$s. Fix a value $2^{i}$. We are trying to “approximate” the hyperbola $\mathcalb{h}\left(2^{i}\right)$ with a sequence whose front roughly coincides with this hyperbola. One way to do so is to consider a sequence $S_{1}$ containing the numbers $\{2^{i}\}^{1},\{2^{i-1}\}^{2},\ldots\{2^{0}\}^{2^{i+1}}$ (not necessarily in this order!).
The intuition for this choice is the following – we might as well round the optimal threshold of the algorithm to the closest (bigger) power of 222, and let $\nabla$ denote this number. Unfortunately, even if we knew that the work of the optimal simulation is $\alpha=2^{i}$, we do not know which value of $\nabla$ to use, so we try all possible choices for this value. One can verify that $S_{i}$ dominates the work hyperbola $\mathcalb{h}\left(2^{i}\right)$ (for all integral points).
Figure 3.3: Plots of three $k$-fronts of TTL sequences generated by the counter strategy. Note that the bars “hide” each other, e.g., the bars of $S_{3}$ (light purple) have $y$-values (8, 4, 4, 2,…). In particular, the region covered by $S_{1}=2,1,1$ is bounded by the blue staircase. Similarly, $S_{2}=4,2,2,1,1,1,1$ is bounded by the red staircase, and $S_{3}=8,4,4,2,2,2,2,1,\ldots,1$ is bounded by the purple staircase.
What is not clear is how to generate an infinite sequence whose prefix is $S_{i}$ if one reads the first $|S_{i}|$ elements of this sequence, for all $i$. As Luby et al. [LSZ93] suggest, ignoring the internal ordering of $S_{i}$, we have that $S_{i+1}=S_{i},S_{i},2^{i+1}$, where $S_{0}=1$. Indeed, every number in $S_{i}$ appears twice in $S_{i+1}$ except for the largest value, which is unique. We have $S_{1}=1,1,2$, $S_{2}=1,1,2,1,1,2,4$, and so on. See Figure 3.3 for an illustration of $k$-fronts generated by the strategy.
Algorithmically, there is a neat way to generate this counter strategy TTL sequence. Let $c$ be an integer counter initialized to 0. To generate the next elements in the sequence, at any point in time, we increase the value of $c$ by one, and add all the numbers $2^{i}$, for $i=0,1,2,3,\ldots$, that divide $c$ into the sequence. Thus, the sequence is
@@
Let $S(i,k)$ be the sum of all the appearances of $2^{i}$ in the first $k$ elements of $T_{k}$. For $2^{i}>2^{j}$ (both appearing in the first $k$ elements), we have $S(i,k)\leq S(j,k)\leq 2S(i,k)$, as one can easily verify.
Let $(1/\beta,\Delta)$ be the profile of alg, see Definition 2.5. The optimal simulation takes in expectation $\mathcal{O}=\Theta(\Delta/\beta)$ time. Let $k$ be the number of elements of $T$ that were used by the simulation before it stopped. Let $B$ be the number of elements in $t_{1},\ldots,t_{k}$, at least as large as $\Delta$. Clearly, $B\sim\mathrm{Geom}(\beta)$ (to be more precise, it is stochastically dominated by it). Thus, we have ${\mathbb{E}}\left[B\right]\leq 1/\beta$. Furthermore, the quantity $\ell=1+\left\lceil{\log_{2}(B\Delta)}\right\rceil$ is an upper bound on the largest index $\ell$ such that $2^{\ell}$ appears in $t_{1},\ldots,t_{k}$. An easy calculation shows that ${\mathbb{E}}\left[\ell\right]=O(1+\log(\Delta/\beta))=O(\log\mathcalb{w}_{\texttt{alg}})$. The total work of the simulation thus is $\sum_{i=0}^{\ell}S(i,k)=O(\ell B\Delta)$, which in expectation is $O(\mathcalb{w}_{\texttt{alg}}\log\mathcalb{w}_{\texttt{alg}})=O(\mathcal{O}\log\mathcal{O})$.
3.2 New restart strategies
In the following, we present and discuss several restarting strategies. We start by introducing randomized TTL strategies, namely random $\zeta$ search and random counter search, in Section 3.3. These strategies randomly generate TTLs whenever a new run starts instead of using a predetermined sequence of TTLs. We prove that both strategies are optimal.
We then introduce two more strategies for which we abandon the “stop/start only” constraint, and allow the use of pausing. The wide search strategy, discussed in Section 3.4, simulates different copies of alg running at different speeds by pausing some runs while allowing others to progress. The cache + X strategy, which is the combination of a caching mechanism and a restarting strategy, maintains a set of paused runs of alg that are “extended” to longer runs instead of simulating alg from scratch at every step. We provide a proof of optimality only for wide search, as the caching mechanism maintains the optimality of the X component of cache + X.
3.3 Parallelizable random search
@@
Intuitively, our goal is to generate a distribution $\mathcal{D}$ such that the $k$-front created by sampling $\mathcal{D}$ $k$ times will approximate the space under the hyperbola $\mathcalb{h}\left(2^{\log k-1}\right)$. These distributions are asymptotically optimal and generate thresholds other than $2^{i}$.
3.3.1 The Zeta 222 distribution
@@
Let $(1/\beta,\Delta)$ be the profile of alg, and let $L=c\cdot\mathcal{O}^{c}$, where $\mathcal{O}=\Theta(\Delta/\beta)$ and $c$ is some sufficiently large constant, let $\mathcal{T}_{L}=t_{1},t_{2},\ldots,t_{L}\sim\zeta(2)$ be the random sequence of TTLs used by the simulation, and let $X\sim\zeta(2)$. A value $t_{i}\in\mathcal{T}_{L}$ is good if $t_{i}\leq L^{2}$. Clearly, the probability that any of the values in $\mathcal{T}_{L}$ is not good is at most $L{\mathbb{P}}\left[X\geq L^{2}\right]=O(1/L)$, since for $i>1$, we have
${\mathbb{P}}\left[X\geq i\right]$ $=\sum_{\ell=i}^{\infty}\frac{6/\pi^{2}}{\ell^{2}}\leq\frac{6}{\pi^{2}}\sum_{\ell=i}^{\infty}\frac{1}{\ell(\ell-1)}$
$=\frac{6}{\pi^{2}}\sum_{\ell=i}^{\infty}\left(\frac{1}{\ell-1}-\frac{1}{\ell}\right)=\frac{6}{\pi^{2}(i-1)}.$
@@
$\tau$ $={\mathbb{E}}\!\left[X\;\middle|\;X\leq L^{2}\right]=\frac{1}{{\mathbb{P}}\left[X\leq L^{2}\right]}\sum_{i=1}^{L^{2}}\frac{6i}{\pi^{2}i^{2}}$
$=O(\log L)=O(\log\mathcal{O}).$
@@ 3.3.2 Random counter search
Consider a generated threshold $t$ as a binary string $b_{1}b_{2}\ldots,b_{k}$, where the numerical value of this string is the value it encodes in base two: $t=\sum_{i=1}^{k}b_{i}2^{k-i}$.
@@
We generate a random string as follows. Let $s_{1}=1$. Next, in each step, the string is finalized with probability 1/21/21 / 2. Otherwise, a random bit is appended to the binary string, where the bit is randomly chosen between 0 or 111 with equal probability. We denote the resulting distribution on the natural numbers BIN. Observe that for $R\sim\textbf{BIN}$, the probability that $R$ length is $k$ (i.e., $\mathcalb{b}\left(R\right)=k$) is $1/2^{k}$, for $k\geq 1$. Also, observe that $R$ has a uniform distribution over all the numbers of the same length.
Note that if we were to “round” every value drawn from the BIN down to the nearest value of the form $2^{i}$, we would get the geometric distribution perfectly mimicking the counter search strategy. We refer to the random strategy using BIN to sample the TTLs as the Random Counter simulation.
@@
[…]
@@
$p$ $={\mathbb{P}}\left[X_{i}=1\right]$
$=\mathop{{\mathbb{P}}}\!\left[\texttt{alg}\text{ successfully runs with {TTL}~}t_{i}\;\middle|\;t_{i}\geq\Delta\right]$
$\quad\quad\cdot{\mathbb{P}}\left[t_{i}\geq\Delta\right]$
@@
The number of values of $T_{L}$ the simulation needs to use until it succeeds is a geometric variable with distribution $\mathrm{Geom}(1/\mathcal{O})$. (Clearly, the probability that the algorithm reads all the values of $T_{L}$ diminishes exponentially, and can be ignored.) Thus, the algorithm in expectation reads $1/p=O(\mathcal{O})$ values of $T_{L}$, and each such value has expectation $O(\log\mathcal{O})$. Thus, conditioned on $\mathcal{G}$, the expected running time of the simulation is $O(\mathcal{O}\log\mathcal{O})$.
3.4 The wide search with pause/resume operations
We now remove the constraint allowing us to only start and terminate runs of alg, and introduce the pause/resume operation, allowing us to freeze a run mid execution, and later resume it with no effect on the performed computations. Note that we still refer to all non-fixed-TTL strategies as “restarting strategies” for the sake of brevity. The wide search strategy uses pause/resume operations to simulate running many copies of the algorithm in parallel, but at different speeds. Specifically, the $i$th copy of alg, denoted by $A_{i}$, is ran at speed $\alpha_{i}=1/i$, for $i=1,\ldots$. So, consider the simulation immediately after it ran $A_{1}$ for $t$ seconds — here $t$ is the current time of the simulation. During that time the $i$th algorithm $A_{i}$ ran $n_{i}(t)=\left\lfloor{t/i}\right\rfloor$ seconds. Naturally, the simulation instantiates $A_{i}$ at time $i$, since $n_{i}(i)=1$, and then runs it for one second. More generally, the simulation allocates a running time of one second to a copy of the algorithm $A_{j}$ at the time $t$, when $n_{j}(t)$ had increased to a new integral value. Then, $A_{j}$ is resumed for one second. This simulation scheme can easily be implemented using a heap, and we skip the low-level details.